OpenStudy (anonymous):

Help!! I'm confused with this one and I don't get it. x(x+6)=9 What's the solution?

OpenStudy (anonymous):

x=3

OpenStudy (lgbasallote):

you use the distributive property to get x^2 + 6x = 9 now you use completing the square

OpenStudy (anonymous):

\[ \begin{align} x(x+6)&=9\\ x^2+6x&=9\\ x^2+6x-9&=0 \end{align} \] Complete the square and solve.

OpenStudy (aravindg):

\[\large x^2+6x-9=0\]

OpenStudy (anonymous):

Yeah. What's the factors?

OpenStudy (anonymous):

\[\huge\ x=-3\]

OpenStudy (anonymous):

Because it can't be (x+3) (x-3)

OpenStudy (anonymous):

It's wrong if you check it

OpenStudy (anonymous):

Have you learned how to complete the square, @chol ?

OpenStudy (anonymous):

I'm confused. If I use (x+3) (x-3) it would be x^2-9 only.

OpenStudy (lgbasallote):

you can NOT use factoring. you use completing the square..so don't transpose 9 :P

OpenStudy (aravindg):

\[(x-3(\sqrt2-1))((x-3(-1-\sqrt2)\]

OpenStudy (anonymous):

Oh what's that?

OpenStudy (aravindg):

=0

OpenStudy (aravindg):

those are the roots

OpenStudy (anonymous):

Radicals? I hate Radicals!!!

OpenStudy (lgbasallote):

lol G...show solutions!! :P

OpenStudy (anonymous):

I don't actually understand the equation o.O

OpenStudy (lgbasallote):

you have x(x-6) = 9 distribute x... x^2 - 6x = 9 do you know how to use completing the square @chol ?

OpenStudy (aravindg):

if u are not good at completing squares just use \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Oh. Ok. I'll try it now on my paper.

OpenStudy (anonymous):

Quadratic formula works but it is good to know how to complete the square.

OpenStudy (lgbasallote):

quadratic formula was derived from completing the square..and completing the square is more useful :) especially in later maths

OpenStudy (anonymous):

Wait, I'm solving. I'll show later what I got

OpenStudy (aravindg):

you guys seem to behave as if i have said "ur method is bad" lols

OpenStudy (anonymous):

Not at all @AravindG, I just wouldn't recommend a person to skip completing the square in favor of the quadratic formula. It's important for her to learn how to complete the square.

OpenStudy (aravindg):

kk

OpenStudy (anonymous):

Oh My! What are you talking about guys? I'm getting more confused!

OpenStudy (lgbasallote):

haha dont mind us :P just solve it hehe

OpenStudy (anonymous):

Hey I got \[-3 \pm 6\sqrt{2}\] Check it please.

OpenStudy (aravindg):

i gave u the answers before

OpenStudy (lgbasallote):

hmm i think it should be \(3\sqrt 2\)

OpenStudy (anonymous):

That is the final answer? The one you showed me earlier?

OpenStudy (aravindg):

yes let me give u the answer @nbouscal and @lgbasallote was speaking of \[\Large (x+3)^2=18\]

OpenStudy (anonymous):

@AravindG : Can you show it to me again? It's not clear. Your answer.

OpenStudy (aravindg):

now u can handle it from there

OpenStudy (aravindg):

\[\Large x+3= \pm \sqrt{18}\] \[\Large x=\pm \sqrt{18}-3\]

OpenStudy (aravindg):

did u understand?

OpenStudy (anonymous):

Yeah. But why use it?

OpenStudy (aravindg):

to get the solution !!

OpenStudy (aravindg):

the solutions are easily got by this method

OpenStudy (aravindg):